OpenStudy (mathmath333):

In four throws with a pair of dices, what is the chance of throwing a double twice?

OpenStudy (mathmath333):

\(\large \color{black}{\begin{align} & \normalsize \text{ In four throws with a pair of dices, what is the chance of}\hspace{.33em}\\~\\ & \normalsize \text{ throwing a double twice? }\hspace{.33em}\\~\\ & a.)\ \dfrac{11}{216} \hspace{.33em}\\~\\ & b.)\ \dfrac{25}{216} \hspace{.33em}\\~\\ & c.)\ \dfrac{35}{216} \hspace{.33em}\\~\\ & d.)\ \dfrac{41}{216} \hspace{.33em}\\~\\ \end{align}}\)

Directrix (directrix):

There are 6 ways to get doubles in one throw of a pair of dice. P(doubles on 1 throw) = 6/36 = 1/6 P(doubles on 1 throw) and P(no doubles on 1 throw) = 1 P(no doubles) = 5/6

Directrix (directrix):

Prob ( 2 double throws out of 4) = C(4,2) * (1/6)^2 * (5/6)^2 =

OpenStudy (mathmath333):



OpenStudy (mathmate):

@mathmath333 doubles could mean double 1, double 2....double 6, so 6/36=1/6. " In four throws with a pair of dices, what is the chance of throwing a double twice? " Assuming the question asks "exactly twice". We know that a throw of 2 dice has 36 outcomes, out of which 6 are doubles. So the probability of throwing a double in one throw of dice is 1/6. Since this probability remains constant, and it's a Bernoulli trial (i.e. either success or failure), and the 4 steps of the experiment are random and independent of each other, we can use the binomial distribution, namely \(P(n,r,p)=nCr~ p^r (1-p)^{n-r}\) where nCr means n choose r = n!/(r!(n-r)!) n=number of steps of experiement=4 r=number of successes=2 p=probability of success at each step. (1/6)

OpenStudy (mathmate):